# Redshift Drift in Relativistic N-Body Simulations

### v1

@@ 4 Results
We here show the results obtained from computing the redshift drift in our simulation and compare to the results expected based on the theoretical prediction of Sec. 3. We computed the redshift drift on the light cone of 10 randomly placed present-time observers. Unless otherwise mentioned, we show results using every one-hundredth particle in the light cone, to keep the data amount manageable.
We start by showing the mean observed redshift drift compared to the expected FLRW background result for all 10 observers in Fig. 1. The mean is calculated in 40 bins. At very low redshifts $z<0.05$ we see a large deviation from the background. This is expected given that the very local environment is likely to be either over- or under-dense compared to the mean. For redshifts $0.40.05<z<$ 0.4, the mean redshift drift fluctuates around the background within approximately $0.5\%$. At the largest redshifts in our light cone, $0.28\lesssim z\lesssim 0.38$, all observers see a mean redshift drift smaller than the expected background result. While this is surprising and would be very interesting if true in general, 10 observers are not enough to draw a definite conclusion, as it may in principle be due to e.g. special features of the observer position. In addition, the deviation is very small and therefore is not likely to have any implications for real-world redshift drift observations. The difference could potentially be due to the integrated Sachs-Wolf effect, i.e. terms that integrate over time derivatives of the Bardeen potentials, arising when considering the contributions of metric fluctuations to the redshift drift, see e.g. [28]. To confirm, one would have to investigate if the effect holds for more observers and higher redhifts.
@@
Next, we focus on one specific observer. In Fig. 3 we show the redshift drift as a function of the redshift for observer 1. We show a hexbin density map [^7] https://matplotlib.org/stable/api/_as_gen/matplotlib.pyplot.hexbin.html [33], which shows the number of particles in hexagonal bins. One can clearly see the expected mean result, but also some surprisingly large scatter, which is of the same order of magnitude as the background signal at low redshifts. This scatter also exhibits some structure. Although the sharp cut at $z\sim 0.15$ is a result of the transition from the full-sky light cone to the partial cone, there seems to be some large scatter in vertical lines at specific redshifts. These result from particles falling into the cluster potential of galaxy clusters in our line of sight and therefore being highly accelerated. We illustrate this in the left panel of Fig. 4, where we show a thin slice through the observer light cone with thickness $\mathrm{d}z=0.01$. We plot the number density of particles and highlight only those particles with $|\delta z-\delta\bar{z}|>0.8\times 10^{-10}$. Zooming in on one of the collections of highlighted particles we can see how they lie along our line of sight, in-front and behind a high over-density. The zoomed-in feature lies at roughly redshift $z=0.23$, where we indeed find one of the mentioned vertical lines in Fig. 3. Since we only show a thin slice of the light cone, we do not see highlighted particles corresponding to each vertical line. Decreasing the limit to $|\delta z-\delta\bar{z}|>0.2\times 10^{-10}$ for highlighted particles in the right panel of Fig. 4, we can see that indeed all of the far outliers lie in over-dense regions/ clusters. We will later see that these particles are the ones with a high peculiar acceleration.
@@
The far outliers in Fig. 3 offer an intriguing possibility of constraining the peculiar acceleration and possibly reconstructing the potential of galaxy clusters. As can be seen from (25) the scatter around the background constrains the Hubble weighted peculiar acceleration field along our line of sight $\mathrm{d}/\mathrm{d}t\,(\mathbb{v}/H)\cdot\mathbb{e}$. We will see in the next subsection that (25) is very accurate. It would be interesting to see how these results could be related to those in [8], where it was first suggested to use the peculiar acceleration to reconstruct the potential of local clusters. To do so, it would be interesting to see how the scatter depends on the mass of the objects studied, e.g. by employing a Halo finder to identify realistic sources.
@@
We further note that since the inhomogeneity-induced fluctuations seen in Fig. 3 are of the same order as the cosmic signal for redshifts up to $z\approx 0.15$, they can lead to a sign change in the redshift drift. Albeit only for the more extreme outliers. The possibility of fluctuations to change the sign of the redshift drift is particularly interesting because the sign of the redshift drift in FLRW spacetimes is a direct indication for ($\delta z>0$) or against ($\delta z<0$) accelerated cosmic expansion. The sign of the redshift drift and its relation to accelerated expansion in an inhomogeneous universe have previously been discussed in e.g. [34; 35; 36; 13] with the conclusion that the sign is still determined by cosmic acceleration, at least for ‘cosmological redshifts’, i.e. with the possible exception of low redshifts where peculiar motion can dominate the signal. Our results support this overall conclusion, although we note that this is the first time a change in the sign of $\delta z$ due to inhomogeneities at redshifts up to $z\approx 0.15$ has been reported. We also note that we will show in the following that the extreme outliers and hence change in the sign of the redshift drift are due to extreme peculiar acceleration. We thus still find that it is the acceleration that determines the sign of the redshift drift, but it is the total local acceleration between emission and observation, not the cosmic acceleration independently.
@@
In Fig. 5 we show HEALPix sky maps of the redshift drift fluctuations at redshifts $z=0.1,0.2,0.3\pm 0.01$ for 3 randomly selected observers. We choose $N_{\mathrm{side}}=256$ for the allsky map at $z\sim 0.1$ and $N_{\mathrm{side}}=512$ for the pencil-beam maps at $z\sim 0.2,0.3$. If there are multiple particles in one cell of the map, we average their redshift drift. In order to be able to see both the outliers and the underlying distribution, the maps are log-scaled between $10^{-10}$ and $10^{-11}$ and linearly scaled between 00 and $10^{-11}$. The distribution follows the large-scale structure and we again find the far outliers in regions of high density/ many particles. The maps for $z=0.3$ have more blue than red points due to the already mentioned shift in the mean with respect to the expected FLRW result.
@@
Figure 5: HEALPix skymaps of the redshift drift fluctuations around the FLRW background prediction, for three different redshift bins $z=0.1,0.2,0.3\pm 0.01$ and three different observers. The map for $z\sim 0.1$ shows the full sky with resolution $N_{\mathrm{side}}=256$, while the maps for $z\sim 0.2,0.3$ show partial circular sections of the sky from a pencil beam light cone with half-opening angle 25° and have a resolution $N_{\mathrm{side}}=512$. These figures are log-scaled between $10^{-10}$ and $10^{-11}$ and linearly scaled between 00 and $10^{-11}$ to allow both visualization of outliers and the underlying distribution.
4.1 Comparison to Theoretical Predictions
@@
The results are shown in Figs. 6,7. In Fig. 6 we again show hexbin density maps, now depicting the redshift drift fluctuations and the individual contributions predicted by (23). The top row shows the redshift drift fluctuations $\delta z-\delta\bar{z}$ in the leftmost panel and the theoretical expectation $\delta z^{v}+\delta z^{\dot{v}}$ in the middle panel; by eye, the two seem to be an almost perfect match. Indeed, when subtracting all contributions from $\delta z$ (top row leftmost panel), we can see that we almost perfectly remove all the scatter. In the bottom row, we show the individual contributions by the acceleration (left) and velocity (right), we can see that the peculiar acceleration is responsible for all the far outliers, while the peculiar velocity produces a much smaller scatter. We have seen earlier in Fig. 4 that all of the far outliers are sitting in over-densities/ cluster. With Fig. 6 we have now confirmed that these are highly accelerated particles that fall into the cluster potentials.
To better quantify how well the predictions from Sect. 3 work, we show normalised histograms for three different redshifts in Fig. 7. We plot the relative difference between the prediction of the theory and the simulated results for $\delta z$ as well as the relative difference, taking only some of the theoretical terms into account. We can see that, when taking all terms into account, Eq. (23) predicts the redshift drift to better than one percent accuracy for almost all particles. If we include only the background and peculiar velocity contribution we get a very narrow distribution close to zero with wide tails, i.e. this simplified theoretical prediction is good for particles with small scatter. If we instead include the background and peculiar acceleration term, we find a wide distribution around zero that rapidly falls of, i.e. we are now predicting particles with large scatter well. This confirms our earlier observations.
@@ 4.2 Angular Power Spectra
For redshifts $z=0.1,0.2,0.3$ we sort those particles with redshift $z\pm 0.01$ into HEALPix maps with $N_{\mathrm{side}}=256$, recording the average value of $\Delta\delta z$, if there are multiple particles in a cell. We now use every tenth particle in the light cone in order to have sufficient statistics and not end up with any empty cells on the HEALPix map. In order to reduce shot noise we create two maps per redshift with half of the particles each, randomly selected, and estimate the auto-correlation spectrum by calculating the cross-spectrum between the two maps. This is known as jackknife resampling (see e.g. [37]). With these maps, we now calculate the spectrum using HEALPix’s function anafast [^8] https://healpix.sourceforge.io/html/fac_anafast.htm. For the two maps at $z=0.2,0.3$ that only have partial sky coverage, we pass a masked map, masking the empty sky regions, divide the spectrum by the sky-fraction $C_{l}/f_{\mathrm{sky}}$ to correct the amplitude, and bin the spectra with $\Delta l\approx 1/f_{\mathrm{sky}}$ in order to reduce statistical noise. For binning the spectra we use the python script bin_llcl.py included with the code PolSpice [^9] https://www2.iap.fr/users/hivon/software/PolSpice [38; 39]. We also calculated the power spectra using PolSpice, but as the results were the same we stuck with the slightly simpler to use anafast.
@@
The resulting power spectra can be seen in Fig. 8. For very small multipole moments $l$ the spectra follow the prediction from linear perturbation theory very well. Depending on the redshift we see a surprisingly sharp increase in power somewhere between $203<l<$ 20. This increase is due to the non-linear signal from the peculiar velocities and acceleration. We ensure the signal is not shot-noise or other type of numerical noise in Appx. B. Such a strong signal from non-linearities for such small $l$ is surprising and has not been seen in earlier simulations of the power spectrum.
@@
As a side-note, we note that the window function has significant effects on the power spectra. This is demonstrated in Fig. 9, where we show the power spectrum for different widths of our window function $\mathrm{d}z=0.001,0.1,0.2$. Here we use all particles in the light cone in order to have sufficient particles to fill the sky map with the narrowest window function, and do this for one observer only due to the numerical cost. One can see that the larger the Tophat, the earlier (smaller $l$) the non-linear signal can be seen. For larger $l$ the spectra seem to somewhat converge independent of the size of the window function.
@@
Power spectra for the redshift drift fluctuations were also calculated from simulations in [13], but included no comparison to predictions from perturbation theory. Using our result from Sect. 3.2 we now make this comparison. In [13], the authors used a fully general relativistic simulation of structure formation in an Einsten-de Sitter (EdS) universe made with the Einstein Toolkit [14; 15]. The initial conditions were set using CLASS, with $h=0.45$ and otherwise standard CLASS parameters. We generate an initial spectrum in the same manner and use the EdS linear growth factor to evolve it to present time for use in (41). The authors of [13] calculated the power spectrum from HEALPix maps with $N_{\mathrm{side}}=32$ for two different redshifts $z=0.1,0.5$ and 10 observers each. Their raytracing algorithm traces to the redshift closest to the two target redshifts. We plot the distribution of redshifts in the individual map pixels for each observer and find that the resulting histogram is best described by a Gaussian with mean $\langle z\rangle=0.108,0.508$ and width $\sigma_{z}=0.002,0.0025$ respectively. We therefore use such Gaussian’s as our window function for calculating the theoretical prediction of the power spectrum. Due to the low number of pixels in the HEALPix maps, this is an approximation, and we cannot expect a perfect match with the simulation results.
@@
The results can be seen in Fig. 10, we show both the prediction using the Gaussian window function (solid grey and black lines) as well as a delta function (dashed grey and black lines) to demonstrate how important the window function is, especially for redshift $z=0.108$. Despite our somewhat crude assumptions, we find a very nice match for low multipole moments $l$ for both redshifts, only slightly overestimating the amplitude. For larger $l$ we find good agreement for $z=0.508$, slightly underestimating the power for the very largest $l$. For redshift $z=0.108$, the theoretical prediction overestimates the amplitude compared to the simulation result at large $l$. There are two possible explanations: i) Our window function was too imprecise, or ii) since these simulations have very little non-linear structure, with the initial spectrum even being cut off for very small scales to avoid numerical noise, it is likely that the simulation just underestimates the power here. This also explains why the non-linear signal we saw in our simulations was not picked up in the simulations of [13]. Overall, the agreement between the simulation results of [13] and our theoretical prediction for the power spectrum seems very good and gives us further confidence that our theoretical result is correct.
The authors of [17] also calculated the redshift drift fluctuation power spectrum in a $\Lambda$CDM universe using N-body simulations. Their simulation was run using the Newtonian N-body code GADGET-4 [18] and had a resolution and cosmological parameters similar to ours. The authors calculated the power spectrum from the peculiar velocities and acceleration on the constant time hypersurfaces of the simulation, i.e. using the background redshift and thus not raytracing to the individual simulation particles. Their power spectra disagree with ours in shape and amplitude [^10] When comparing amplitudes, one has to account for the different normalizations employed by [17] and us. See the discussion in Appx. A. This difference was taken into account in the comparison.. The amplitudes in [17] also disagree with the perturbation theory results by [28], but unfortunately the agreement is too large to be fixed by the disagreement we found with [28] in this paper, see Appx. A. Since our simulation results agree very well with the predictions from perturbation theory and we have demonstrated that calculating the redshift drift from the peculiar velocity and acceleration is an excellent approximation, it seems most likely that the approximation of using the background redshift and not raytracing causes the discrepancies between the power spectra.
@@
Figure 10: Comparison between predictions for the angular power spectrum of redshift drift fluctuations from perturbation theory as calculated in this paper and simulation results by [13].

### v2

@@ 4 Results
We here show the results obtained from computing the redshift drift in our simulation and compare to the results expected based on the predictions of Sec. 3. We computed the redshift drift on the light cone of 10 randomly placed present-time observers. Unless otherwise mentioned, we show results using every one-hundredth particle in the light cone, to keep the data amount manageable.
We start by showing the mean observed redshift drift compared to the expected FLRW background result for all 10 observers in Fig. 1. The mean is calculated in 40 bins. At very low redshifts $z<0.05$ we see a large deviation from the background. This is expected given that the very local environment is likely to be either over- or under-dense compared to the mean. For redshifts $0.05<z<0.4$, the mean redshift drift fluctuates around the background within approximately $0.5\%$. At the largest redshifts in our light cone, $0.28\lesssim z\lesssim 0.38$, all observers see a mean redshift drift smaller than the expected background result. While this is surprising and would be very interesting if true in general, 10 observers are not enough to draw a definite conclusion, as it may in principle be due to e.g. special features of the observer position. In addition, the deviation is very small and therefore is not likely to have any implications for real-world redshift drift observations. The difference could potentially be due to the integrated Sachs-Wolf (ISW) effect, i.e. terms that integrate over time derivatives of the Bardeen potentials, arising when considering the contributions of metric fluctuations to the redshift drift, see e.g. [Bessa2023]. To confirm, one would have to investigate if the effect holds for more observers and higher redhifts.
@@
Next, we focus on one specific observer. In Fig. 3 we show the redshift drift as a function of the redshift for observer 1. We show a hexbin density map [^7] https://matplotlib.org/stable/api/_as_gen/matplotlib.pyplot.hexbin.html [matplotlib], which shows the number of particles in hexagonal bins. One can clearly see the expected mean result, but also some surprisingly large scatter, which is of the same order of magnitude as the background signal at low redshifts. This scatter also exhibits some structure. Although the sharp cut at $z\sim 0.15$ is a result of the transition from the full-sky light cone to the partial cone, there seems to be some large scatter in vertical lines at specific redshifts. These result from particles falling into the cluster potential of galaxy clusters in our line of sight and therefore being highly accelerated. We illustrate this in the left panel of Fig. 4, where we show a thin slice through the observer light cone with thickness $\mathrm{d}z=0.01$. We plot the number density of particles and highlight only those particles with $|\delta z-\delta\bar{z}|>0.8\times 10^{-10}$. Zooming in on one of the collections of highlighted particles we can see how they lie along our line of sight, in front of and behind a high over-density. The zoomed-in feature lies at roughly redshift $z=0.23$, where we indeed find one of the mentioned vertical lines in Fig. 3. Since we only show a thin slice of the light cone, we do not see highlighted particles corresponding to each vertical line. Decreasing the limit to $|\delta z-\delta\bar{z}|>0.2\times 10^{-10}$ for highlighted particles in the right panel of Fig. 4, we can see that indeed all of the far outliers lie in over-dense regions/ clusters. We will later see that these particles are the ones with a high peculiar acceleration.
@@
The far outliers in Fig. 3 offer an intriguing possibility of constraining the peculiar acceleration and possibly reconstructing the potential of galaxy clusters. As can be seen from (25) the scatter around the background constrains the Hubble weighted peculiar acceleration field along our line of sight $\mathrm{d}/\mathrm{d}t\,(\mathbb{v}/H)\cdot\mathbb{e}$. We will see in the next subsection that (25) is very accurate. It would be interesting to see how these results could be related to those in [Amendola2008], where it was first suggested to use the peculiar acceleration to reconstruct the potential of local clusters. To do so, it would be interesting to see how the scatter depends on the mass of the objects studied, e.g. by employing a Halo finder to identify realistic sources.
@@
We further note that since the inhomogeneity-induced fluctuations seen in Fig. 3 are of the same order as the cosmic signal for redshifts up to $z\approx 0.15$, they can lead to a sign change in the redshift drift, albeit only for the more extreme outliers. The possibility of fluctuations to change the sign of the redshift drift is particularly interesting because the sign of the redshift drift in FLRW spacetimes is a direct indication for ($\delta z>0$) or against ($\delta z<0$) accelerated cosmic expansion. The sign of the redshift drift and its relation to accelerated expansion in an inhomogeneous universe have previously been discussed in e.g. [Koksbang2019; Koksbang2020; Heinesen2021b; Koksbang2024] with the conclusion that the sign is still determined by cosmic acceleration, at least for ‘cosmological redshifts’, i.e. with the possible exception of low redshifts where peculiar motion can dominate the signal. Our results support this overall conclusion, although we note that this is the first time a change in the sign of $\delta z$ due to inhomogeneities at redshifts up to $z\approx 0.15$ has been reported. We also note that we will show in the following that the extreme outliers and hence change in the sign of the redshift drift are due to extreme peculiar acceleration. We thus still find that it is the acceleration that determines the sign of the redshift drift, but it is the total local acceleration between emission and observation, not the cosmic acceleration independently.
@@
In Fig. 5 we show HEALPix sky maps of the redshift drift fluctuations at redshifts $z=0.1,0.2,0.3\pm 0.01$ for 3 randomly selected observers. We choose $N_{\mathrm{side}}=256$ for the allsky map at $z\sim 0.1$ and $N_{\mathrm{side}}=512$ for the pencil-beam maps at $z\sim 0.2,0.3$. If there are multiple particles in one cell of the map, we average their redshift drift. In order to be able to see both the outliers and the underlying distribution, the maps are log-scaled between $10^{-10}$ and $10^{-11}$ and linearly scaled between 0 and $10^{-11}$. The distribution follows the large scale structure and we again find the far outliers in regions of high density/ many particles. The maps for $z=0.3$ have more blue than red points due to the already mentioned shift in the mean with respect to the expected FLRW result.
@@
Figure 5: HEALPix skymaps of the redshift drift fluctuations around the FLRW background prediction, for three different redshift bins $z=0.1,0.2,0.3\pm 0.01$ and three different observers. The map for $z\sim 0.1$ shows the full sky with resolution $N_{\mathrm{side}}=256$, while the maps for $z\sim 0.2,0.3$ show partial circular sections of the sky from a pencil beam light cone with half-opening angle 25° and have a resolution $N_{\mathrm{side}}=512$. These figures are log-scaled between $10^{-10}$ and $10^{-11}$ and linearly scaled between 0 and $10^{-11}$ to allow both visualization of outliers and the underlying distribution.
4.1 Comparison to Analytical Predictions
@@
The results are shown in Figs. 6,7. In Fig. 6 we again show hexbin density maps, now depicting the redshift drift fluctuations and the individual contributions predicted by (23). The top row shows the redshift drift fluctuations $\delta z-\delta\bar{z}$ in the leftmost panel and the analytical expectation $\delta z^{v}+\delta z^{\dot{v}}$ in the middle panel; by eye, the two seem to be an almost perfect match. Indeed, when subtracting all contributions from $\delta z$ (top row leftmost panel), we can see that we almost perfectly remove all the scatter. In the bottom row, we show the individual contributions by the acceleration (left) and velocity (right). We see that the peculiar acceleration is responsible for all the far outliers, while the peculiar velocity produces a much smaller scatter. We have seen earlier in Fig. 4 that all of the far outliers are sitting in over-densities/ cluster. With Fig. 6 we have now confirmed that these are highly accelerated particles that fall into the cluster potentials.
To better quantify how well the predictions from Sect. 3 work, we show normalised histograms for three different redshifts in Fig. 7. We plot the relative difference between the prediction of the analytical and the simulated results for $\delta z$ as well as the relative difference, taking only some of the analytically derived terms into account. We can see that, when taking all terms into account, Eq. (23) predicts the redshift drift to better than one percent accuracy for almost all particles. If we include only the background and peculiar velocity contributions we get a very narrow distribution close to zero with wide tails, i.e. this simplified theoretical prediction is good for particles with small scatter. If we instead include the background and peculiar acceleration term, we find a wide distribution around zero that rapidly falls off, i.e. we are now predicting particles with large scatter well. This confirms our earlier observations.
@@ 4.2 Angular Power Spectra
For redshifts $z=0.1,0.2,0.3$ we sort those particles with redshift $z\pm 0.01$ into HEALPix maps with $N_{\mathrm{side}}=256$, recording the average value of $\Delta\delta z$, if there are multiple particles in a cell. We now use every tenth particle in the light cone in order to have sufficient statistics and not end up with any empty cells on the HEALPix map. In order to reduce shot noise we create two maps per redshift with half of the particles each, with particles being split randomly into the two maps. We then estimate the auto-correlation spectrum by calculating the cross-spectrum between the two maps. This is known as jackknife resampling (see e.g. [Lepori2021]). With these maps, we now calculate the spectrum using HEALPix’s function anafast [^8] https://healpix.sourceforge.io/html/fac_anafast.htm. For the two maps at $z=0.2,0.3$, which only have partial sky coverage, we pass a masked map, masking the empty sky regions, divide the spectrum by the sky-fraction $C_{l}/f_{\mathrm{sky}}$ to correct the amplitude, and bin the spectra with $\Delta l\approx 1/f_{\mathrm{sky}}$ in order to reduce statistical noise. For binning the spectra we use the python script bin_llcl.py included with the code PolSpice [^9] https://www2.iap.fr/users/hivon/software/PolSpice [Chon2004; Szapudi2001]. As a cross-check, we also calculated the power spectra using PolSpice, confirming that results we obtained with anafast and PolSpice agree.
@@
The resulting power spectra can be seen in Fig. 8. For very small multipole moments $l$ the spectra follow the prediction from linear perturbation theory very well. Depending on the redshift we see a surprisingly sharp increase in power somewhere between $3<l<20$. This increase is due to the non-linear signal from the peculiar velocities and acceleration. We ensure the signal is not shot-noise or other type of numerical noise in Appx. B. Such a strong signal from non-linearities for such small $l$ is surprising and has not been seen in earlier simulations of the redshift drift power spectrum.
@@
As a side-note, we note that the window function has significant effects on the power spectra. This is demonstrated in Fig. 9, where we show the power spectrum for different widths of our window function $\mathrm{d}z=0.001,0.1,0.2$. Here, we use all particles in the light cone in order to have sufficient particles to fill the sky map with the narrowest window function, and do this for one observer only due to the numerical cost. One can see that the larger the Tophat, the earlier (smaller $l$) the non-linear signal can be seen. For larger $l$ the spectra seem to somewhat converge independent of the size of the window function.
@@
Power spectra for the redshift drift fluctuations were also calculated from simulations in [Koksbang2024], but included no comparison to predictions from perturbation theory. Using our result from Sect. 3.2 we now make this comparison. In [Koksbang2024], the authors used a fully general relativistic simulation of structure formation in an Einsten-de Sitter (EdS) universe made with the Einstein Toolkit [Macpherson2017; Macpherson2019]. The initial conditions were set using CLASS, with $h=0.45$ and otherwise standard CLASS parameters. We generate an initial spectrum in the same manner and use the EdS linear growth factor to evolve it to present time for use in (41). The authors of [Koksbang2024] calculated the power spectrum from HEALPix maps with $N_{\mathrm{side}}=32$ for two different redshifts $z=0.1,0.5$ and 10 observers each. Their raytracing algorithm traces to the redshift closest to the two target redshifts. We plot the distribution of redshifts in the individual map pixels for each observer and find that the resulting histogram is best described by a Gaussian with mean $\langle z\rangle=0.108,0.508$ and width $\sigma_{z}=0.002,0.0025$ respectively. We therefore use such Gaussian’s as our window function for calculating the linear prediction of the power spectrum. Due to the low number of pixels in the HEALPix maps, this is an approximation, and we cannot expect a perfect match with the simulation results.
@@
The results can be seen in Fig. 10 where we show both the prediction using the Gaussian window function (solid grey and black lines) as well as a delta function (dashed grey and black lines) to demonstrate how important the window function is, especially for redshift $z=0.108$. Despite our somewhat crude assumptions, we find a very nice match for low multipole moments $l$ for both redshifts, only slightly overestimating the amplitude. For larger $l$ we find good agreement for $z=0.508$, slightly underestimating the power for the very largest $l$. For redshift $z=0.108$, the linear prediction overestimates the amplitude compared to the simulation result at large $l$. There are two possible explanations: i) Our window function was too imprecise, or ii) since these simulations have very little non-linear structure, with the initial spectrum even being cut off for very small scales to avoid numerical noise, it is likely that the simulation just underestimates the power here. This also explains why the non-linear signal we saw in our simulations was not picked up in the simulations of [Koksbang2024]. Overall, the agreement between the simulation results of [Koksbang2024] and our linear predictions for the power spectrum seems reasonable and gives us further confidence that our linear result is correct.
[…]
@@
Figure 10: Comparison between predictions for the angular power spectrum of redshift drift fluctuations from perturbation theory as calculated in this paper and simulation results by [Koksbang2024].
5 Implication for Observations
Observations of the redshift drift with the SKA will be obtained by observing $\sim 10^{7}$ galaxies per redshift bin and stacking their signal, allowing a detection in the redshift range $0.2\lesssim z\lesssim 1$ [Kloeckner2015]. More galaxies will be observed at lower redshifts, allowing a stronger detection at the lower end of this range. Due to the necessity of stacking many galaxies, such an experiment will constrain the mean/background signal of the redshift drift, averaging out fluctuations. Our results confirm that there is no significant bias to the mean from peculiar velocities or relativistic effects up to redshifts $z\approx 0.38$. When comparing the mean redshift drift of our simulations with the background value, we find only a very minor deviation between the two across all ten considered observers. This strongly indicates that there is no significant observer-dependent bias from the large scale cosmic neighbourhood of the observer.
Observations of the redshift drift with the ELT will be obtained at much higher redshifts ($2\lesssim z\lesssim 5$) [Liske2008] by studying quasar absorption lines. While our current simulations do not extend to such high redshifts, we have for the first time presented a numerical framework that can in principle accurately calculate the redshift drift signal including relativistic corrections (such as gravitational time dilation and metric perturbations) and light cone effects. This setup can be used to study the high redshift signal and check for potential biases e.g. due to ISW-like terms. Earlier estimates indicated that peculiar velocities will not bias the observed signal [Liske2008; Cooke2020]. Explicitly demonstrating this with relativistic, non-linear simulations is an important step for ensuring reliable observational strategies and for building confidence in future detections. Due to the high computational cost, we leave a detailed high-redshift study to future work. Nonetheless, we have here already shown that at least at low redshifts, accounting for fluctuations from peculiar velocities only, and neglecting any relativistic corrections, allows us to reconstruct the redshift drift signal to better than one percent, independent of redshift. We expect that this would also be the case at high redshifts, except for potential integrated effects that grow with redshift, such as the ISW effect.
We also note that the strong non-linear signals we have described, while not observable with the current telescopes, could provide valuable insight into the gravitational potentials inside clusters and thereby provide strong constraints on modified theories of gravity. The prospect of detecting such a signal could provide motivation for future telescope missions, such as suggested in e.g. [Eikenberry2019].
Lastly we note that although the peculiar motion of sources has only small impact on expected redshift drift measurements with SKA and ELT, the effects may be more significant for alternative methods. This may for instance be the case for the proposed redshift difference which utilizes strong gravitational lensing to detect changes in the redshift due to different travel times [Wang2022; Wang2023].